OpenStudy (anonymous):

you flip three coins what is the probability that you get exactly 3 heads , given that you get at least 1 head ?

OpenStudy (anonymous):

P(A|B) = P(AB)/P(B) A = You get exactly 3 heads. B = You get at least 1 head.

OpenStudy (zarkon):

\[\begin{array}{c|cccc} x & 0 & 1 & 2 & 3 \\\hline P(X=x) & \frac{1}{8} & \frac{3}{8} & \frac{3}{8} & \frac{1}{8} \end{array}\]

OpenStudy (anonymous):

so its 1/8 ?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

use mtbender74 formula with the distribution I gave you

OpenStudy (zarkon):

noting that AB in this context means A and B.

OpenStudy (anonymous):

Also noting that P(B) is P(X=1) + P(X=2) + P(X=3)...

OpenStudy (anonymous):

1/7 :)

OpenStudy (zarkon):

yes